# Modular Compilation for Quantum Chiplet Architectures

### v3

@@ 3.2.1. Stratification Stage
2:$q_{0},q_{1}\leftarrow\textsc{Qubits}(g)$
3:$s_{0},s_{1}\leftarrow L[q_{0}],L[q_{1}]$
4:Increment $C$ if $s_{0}\neq s_{1}$EndFor
[…]
7:$s_{0}\leftarrow\textsc{RandChoice}(\{1,\dots,N_{C}\})$
8:$q_{0}\leftarrow\textsc{RandChoice}(L^{\prime}[s_{0}])$
9:$s_{1}\leftarrow\textsc{RandChoice}(\{1,\dots,N_{C}\}-\{s_{0}\})$
10:$q_{1}\leftarrow\textsc{RandChoice}(L^{\prime}[s_{1}])$LCommentSwap $q_{0}$ and $q_{1}$
11:$L^{\prime}[s_{0}]\leftarrow(L^{\prime}[s_{0}]\cup\{q_{1}\})-\{q_{0}\}$
12:$L^{\prime}[s_{1}]\leftarrow(L^{\prime}[s_{1}]\cup\{q_{0}\})-\{q_{1}\}$EndFor
[…]
14:$L\leftarrow$ Initial Layout $\triangleright$ Current Solution
15:$R\leftarrow L$triangleright Best Solution
[…]
17:$D\leftarrow\textsc{Ceil}(T\times\alpha)$triangleright Number of changes to make
18:$L^{\prime}\leftarrow\textsc{Kick}(D,N_{C},N_{Q},L)$
19:$\Delta\leftarrow\textsc{Cost}(Circ,L^{\prime})-\textsc{Cost}(Circ,L)$
20:$k\leftarrow\textsc{Rand}(0,1)$triangleright Random number in $[0,1)\subset\mathbb{R}$
21:$L\leftarrow L^{\prime}$ if $(\Delta<0)$ or $(k>\exp(\Delta/T))$
22:$R\leftarrow L^{\prime}$ if $(\textsc{Cost}(Circ,L^{\prime})<\textsc{Cost}(Circ,R))$
[…]
24:\$R$EndProcedure
[…]
Qubit-to-subcircuit mapping produces a stratified quantum circuit, where logical qubits (numbers in Figure 3(b)) are grouped into subcircuits (lowercase letters). An optimal qubit-to-subcircuit mapping minimizes the number of connections between subcircuits, thereby reducing the number of inter-chiplet SWAPs needed to transfer data between chiplets during execution. To perform this task, we investigate two methods: Simulated Annealing and Spatio-Temporal Aware Qubit Allocation (STA) (ovide2024sta).
Simulated annealing is an efficient stochastic optimization technique shown to work well on NP-complete problems, such as place-and-route in VLSI algorithms (annealing1989). Simulated annealing is an iterative method to optimize a state against a given cost function. In each iteration, the state is “kicked,” or randomly perturbed, and the new state is greedily accepted if there is a reduction in cost. However, the algorithm may randomly also accept worse solutions to escape from local minima wells. The runtime of simulated annealing is statically determined using constant tuning parameters, thus the overall complexity is determined by the complexity of the cost and kick functions, which are typically polynomial. Additionally, since each run of a simulated annealing is randomized and independent, it is possible to run multiple trials in parallel with different initial random seeds, then pick the best result. This improves the chances of finding a (near-)optimal solution. Our method is shown in Algorithm 1. We start with an initial qubit-to-subcircuit mapping, which we minimize against a cost function, defined on line 1 as the number of inter-chiplet gates given a certain layout. By minimizing the number of inter-chiplet gates, we implicitly reduce the inter-chiplet connectivity. The kick function then swaps random qubits in the layout, ensuring that each SWAP is between qubits in different subcircuits. In doing so, we hope to move strongly interacting qubits close together, minimizing the number of cross-chiplet gates. As input, our algorithm takes the input circuit $Circ$, the chiplet count $N_{C}$, the chiplet size $N_{Q}$, the starting temperature $T_{0}$, the number of changes to do per degree of temperature $\alpha$, and the cooling factor $\beta$. The last three are constant tuning parameters across runs.
The STA algorithm is a qubit allocation algorithm designed for trapped-ion quantum systems. These devices often feature modular designs, where one device is composed of “traps” arranged in a (typically) line or ring topology, where each trap contains multiple qubits (ions). Thus, it seems possible to adapt qubit allocation algorithms designed for trapped-ion systems to support modular transmon systems as well. Specifically, one component of STA, denoted as Algorithm 2 in Ovide et al. (ovide2024sta), assigns qubits to ion traps, which we can leverage to perform qubit-to-subcircuit mapping and generate an initial subcircuit-to-chiplet mapping.
Chiplet Allocation
Algorithm 2 Inter-Chiplet Routing
[…]
8:$V_{virt},V_{phys}\leftarrow\textsc{ValidGates}(Q,L)$
[…]
12:$L^{\prime}\leftarrow\textsc{Swap}(\textsc{Copy}(L),S)$
[…]
14:$q_{0},q_{1}\leftarrow\textsc{Qubits}(g)$
15:$s_{0},s_{1}\leftarrow L[q_{0}],L[q_{1}]$
[…]
18:$D\leftarrow\textsc{CircuitToDAG}(Circ)$
19:$E\leftarrow$ All-pairs distances from chiplet topology $T$
20:$Q\leftarrow\textsc{FirstLayer}(DAG)$
21:$R\leftarrow$ empty list
[…]
23:$\mathcal{S}\leftarrow\textnormal{{FindSwaps}}(Q,E)$triangleright SWAP Candidates
24:$S\leftarrow\operatorname*{arg\,min}_{s\in\mathcal{S}}\textsc{Heuristic}(s,R,Circ)$
25:$\textsc{Swap}(L,S)$
26:$\textsc{Step}(Q,R,L,D)$EndWhile
27:$D^{\prime}\leftarrow$ Empty DAG
28:Append $g$ to $D^{\prime}$ for each $g\in{R}$triangleright $R$ is ordered
[…]
Chiplet allocation describes the process of transforming a stratified circuit to be compatible with the inter-chiplet topology of a modular device. In Figure 2, we decompose chiplet allocation into subcircuit-to-chiplet mapping and inter-chiplet routing, analogous to how qubit allocation can be decomposed into qubit layout and routing. Accordingly, subcircuit-to-chiplet mapping (Figure 3(c)), defines an initial placement of logical subcircuits (lowercase letters) to physical chiplets (uppercase letters). Subsequently, inter-chiplet routing (Figure 3(d)) inserts inter-chiplet SWAPs to account for limited cross-chiplet gate connectivity.
In our design, we use our qubit-to-subcircuit mapping to provide an initial subcircuit-to-chiplet mapping. Then, we extend and modify the SABRE algorithm (sabre_swap) to apply to chiplets, placing iterations of inter-chiplet routing alongside randomized qubit permutations to inform a final subcircuit-to-chiplet mapping. We did explore a design that extended other components of STA to perform subcircuit-to-chiplet mapping on transmon chiplet systems, but did not produce circuits meaningfully better than this iterative approach and increased compilation time, hence it was abandoned.
Inter-chiplet routing seeks to consolidate the qubits of (non-SWAP) inter-chiplet gates to a single chiplet by inserting inter-chiplet SWAPs into the stratified circuit. For a single two-qubit, inter-chiplet gate, we represent inter-chiplet routing as a shortest-path problem, where the two qubits of the gate correspond to chiplet endpoints in the unweighted inter-chiplet connectivity graph. From the shortest path, we can derive an optimal sequence of SWAPs for routing, where the unweighted path length, or chiplet distance, indicates the length of the SWAP chain. However, in the context of a stratified circuit, composed of multiple inter-chiplet gates, it becomes possible to interleave SWAP chains (i.e., efficiently reorder their execution) to reduce the depth and/or total number of inter-chiplet SWAPs. Like qubit routing, we would expect it to be impractical to derive an optimal solution for this more complex picture of inter-chiplet routing at compile-time. Thus, we employ a similar SWAP selection scheme as SABRE (sabre_swap), where inserted SWAPs are chosen from a pool of SWAP candidates according to a cost heuristic. For our inter-chiplet routing algorithm, we categorize SWAPs from our candidate pool into three tiers, based on a concept of the “efficiency” of each SWAP, as shown in Algorithm 2.
An inter-chiplet SWAP is considered symbiotic if it directly affects two inter-chiplet gates in a “beneficial” manner. In other words, each qubit of a symbiotic SWAP corresponds to a different inter-chiplet gate, where the insertion of the SWAP reduces the chiplet distance score of both gates. In theory, this outcome represents the most efficient possible insertion for a single SWAP, and thus should be the highest-priority case for our greedy inter-chiplet routing cost heuristic. Next, commensalistic SWAPs are those that do not “harm,” or increase that chiplet distance, of any directly impacted gates. These SWAPs typically occur when an idle qubit can be used as an ancilla qubit to facilitate the routing of an inter-chiplet gate without affecting other inter-chiplet gates. In certain device topologies, a SWAP acting on the qubits of two inter-chiplet gates can also exhibit commensalistic behavior. Finally, parasitic SWAPs act on the qubits of two inter-chiplet gates, where one gate benefits while the other is harmed. Since this variety of SWAPs is actively counterproductive towards a gate, it should be, and likely will be, avoided in most circumstances. However, given a particularly onerous circuit, it may be necessary in such rare instances to break out of a deadlock. For example, clearing one gate may free physical qubits to help route others.
3.2.2. Elaboration Stage
(a) An intra-chiplet topology
(b) A $3$-qubit subcircuit post-stratification
(c) Subcircuit after intra-chiplet layout
(d) Inter-chiplet SWAP lowering
(e) Intra-chiplet routing
(f) Basis translation and optimization
Figure 4. Step-by-step example of intra-chiplet compilation in the Elaboration stage. Gates that cannot be implemented on hardware are shown in red.
Following the stratification stage, each chiplet can be compiled almost completely independently from one another. More precisely, the compilation process within each chiplet is largely identical to that of a monolithic quantum architecture. The only exception is the presence of inter-chiplet SWAPs, which must be lowered from the chiplet level to the qubit level. Once this assignment is made, these SWAPs must remain fixed and immutable to subsequent compilation stages. Despite this overhead, restricting the task of compilation to a narrower domain of qubits facilitates both parallelization (since each subcircuit can be processed in parallel) and problem size reduction (since there are fewer qubits to consider) of the most computationally intensive compilation stages, namely routing and optimization. To accomplish these tasks, we divide the elaboration stage into two substages: qubit allocation and parallel basis translation and optimization. Figure 4 illustrates this with an example of elaboration for a single chiplet (Figure 4(a)) and its corresponding subcircuit (Figure 4(b)).
Qubit Allocation
We employ a modified version of SABRE to perform qubit allocation in the elaboration stage. Specifically, we parallelize the SABRE routing algorithm and add constraints to maintain correctness when applied to a stratified quantum circuit with inter-chiplet gates. We also explored a layout pass that used portions of the STA algorithm, but the marginal benefits to solution quality were not worth the 56×5-6$\times$5 - 6 × slower elaboration time.
The placement of inter-chiplet gates is a non-parallelizable task (due to cross-chiplet interactions) that strongly influences the layout and routing passes, which could otherwise be fully parallelized across chiplets. To account for this, we introduce a serial stage (Figure 4(d)) between the parallel layout and routing passes that greedily places gates on the nearest valid inter-chiplet edge. A greedy policy was chosen because inter-chiplet gates have more restrictions on valid placements compared to intra-chiplet gates, while also experiencing substantially higher error rates (smith_2022_chiplets). Once the positions of the inter-chiplet gates are settled, they must remain fixed, imposing an additional constraint on qubit routing. To promote this behavior in SABRE, we extend its cost heuristic to weigh the device topology by fidelity. This causes higher error links to assume higher costs, discouraging SABRE from moving the high-error inter-chiplet gates. As a result, SEQC achieves better solutions in the common case. However, there is nothing to prevent SABRE from incorrectly modifying any inter-chiplet SWAPs. Thus, we need to impose hard restrictions to ensure correctness. We address this by establishing two disjoint gate sets. One gate set operates strictly within a chiplet. This is the traditional gate set for a monolithic processor. The other gate set is the only one permitted to operate across chiplets, and specifically only implements inter-chiplet SWAPs. By restricting each gate set to its own disjoint part of the overall device topology, we ensure correctness across all cases.
Parallel Translation and Optimization
The remaining compilation stages of basis translation and circuit optimization can be easily parallelized at the chiplet granularity (Figure 4(f)). Basis translation, in particular, is embarrassingly parallel: gates can be translated independently from each other. Meanwhile, as discussed in §2, optimization algorithms already employ partitioning and parallelization to create tractable subproblems from intractably large circuits, facilitating a smaller problem space, lower memory usage, and faster performance. In modular quantum architectures, we physically motivate partitions through the natural chiplet boundaries, enabling us to reap the same resource and performance benefits of partitioning without compromising on solution quality.
4. Methodology
We evaluate quantum compilers based on the quality of their solutions, their speed of compilation, and their ability to scale to large circuits/devices. To experimentally profile these characteristics, we benchmark a compiler on a suite of quantum circuits, detailed in §4.1, targeting a mock chiplet-based quantum backend, as specified in §4.2. In total, we evaluate four compilers—a reference compiler, a peephole-augmented compiler, and two versions of SEQC—on simulated backends ranging from 1 to 100 chiplets. All compilers are implemented in the Qiskit SDK (qiskit2024) v1.2.41.2.41.2.41.2.4 and Python $3.12.6$, and timing results are taken on a machine with a $128$-core Ampere Altra Max ARM64 processor and $256$ GB of DDR4 RAM. For our reference compiler, we use the one built into Qiskit, parameterized with optimization_level=3. Because Qiskit is not natively chiplet-aware, we run these experiments on a backend with universal inter-chiplet links (§4.2). To construct the peephole compiler we augment the reference Qiskit compiler by inserting our peephole pass between its routing and basis_translation stages, thus making it chiplet-aware. Our two SEQC versions differ in the technique they use for qubit-to-subcircuit mapping. Our first SEQC implementation leverages simulated annealing, initialized with a starting temperature $T_{0}=200$, a rate of change of $0.005T$ per round, $\frac{T}{50}$ permutations per round, and a total of $10$ trials per CPU core. Meanwhile, our second SEQC implementation leverages STA for subcircuit-to-qubit mapping and subcircuit-to-chiplet mapping.
We aggregate results in a manner similar to the SPEC family of benchmarks (spec2017), by taking the geometric mean of the relative performance ratio against a baseline technique across a benchmark suite of quantum circuits. The resultant score is then derived for each of the evaluated metrics on multiple backends of varying numbers of chiplets.
4.1. Benchmark Suite
@@
[…]
Greenberger–Horne–Zeiling (GHZ). Entanglement is one of the fundamental properties of quantum states and a major source of quantum advantage. The GHZ state represents a maximally entangled $n$-qubit state of $\frac{1}{\sqrt{2}}\left(\ket{|}0>^{\otimes n}+\ket{|}1>^{\otimes n}\right)$. The algorithmic circuit structure is essentially a single chain of CNOT gates (Figure 5(c)) From a compilation perspective, this represents a pure exercise in qubit allocation, because the logical circuit is entirely serial. An $n$-qubit GHZ circuit that fully utilizes a $k$-chiplet device should necessitate $\lfloor\frac{n}{k}\rfloor-1$ logical CNOT operations to extend between chiplets. In turn, this will necessitate at least $2\left(\lfloor\frac{n}{k}\rfloor-1\right)$ inter-chiplet SWAPs to transfer qubit in and out of a chiplet.
@@ 4.2. Simulated Quantum Device Specifications
We perform our experiments on mock backends that are generated to conform to the chiplet architecture proposed in Smith et al. (smith_2022_chiplets). The hardware specifications of the qubits and intra-chiplet connections are sourced from Acharya et al. (backend_specs). Table 1 details the full specifications. To generate the expected fidelity of the inter-chiplet SWAPs, we perform a simulated random benchmarking trial and calculate the error rate of a SWAP gate with these backend specifications. Following Smith et al. (smith_2022_chiplets), we model inter-chiplet SWAPs on this backend with 4×4$\times$4 × the intra-chiplet error rate. Similarly, we approximate the duration of the inter-chiplet SWAP gate as 4×4$\times$4 × the duration of an intra-chiplet SWAP.
@@
CZ (intra-chiplet) (backend_specs) $34.0$ $0.605\%$
CZ (inter-chiplet) $126.0$ $2.420\%$
SWAP (inter-chiplet) $702.4$ $10.23\%$
@@
Table 1. Specifications of Simulated Chiplet Backend.
The simulated device has a grid lattice topology of symmetric chiplet modules, where each chiplet features a heavy-hexagon qubit topology (heavy_hex) (Figure 6). The intra-chiplet topology is constructed such that the heavy-hex structure is preserved for the global device topology (ignoring chiplet division). The tessellation requirements of each chiplet restrict valid chiplet sizes to multiples of $10$; however, for our experiments, we limit the size of each chiplet to the minimum of $10$ qubits due to hardware limitations. Finally, note that for each valid chiplet count, there may exist multiple valid chiplet topologies. In this work, we generate our backends to use the “most-square” topology for a given chiplet count. For example, a $12$-chiplet machine would be generated with a grid of $3\times 4$ chiplets.
Figure 6. Construction of a grid inter-chiplet lattice from a heavy-hexagon qubit lattice, where the heavy-hex lattice is maintained for each chiplet.
@@ 5. Experimental Results
We present results for the Qiskit reference compiler, the peephole-augmented Qiskit (PA-Qiskit) compiler, the SEQC compiler using simulated annealing (SEQC-AN), and the SEQC compiler using STA (SEQC-STA). The Qiskit-based and SEQC compilers prioritize different heuristics in compilation. We examine how well each compiler is able to minimize its preferred heuristics, and how well that translates to practical solution quality. Finally, we analyze compilation time, specifically the recurring and non-recurring compilation times for repeating compilations of a single application.
Number of inter-chiplet gates
Figure 7. Relative number of inter-chiplet gates for SEQC normalized to the Qiskit baseline (lower is better).
@@
The primary heuristic we minimize in SEQC is the number of inter-chiplet gates in a quantum circuit. This is because we expect that the inter-chiplet gates of a chiplet architecture contribute the most to error and execution time relative to other gates (at least by a factor of 4×4$\times$4 × (smith_2022_chiplets)). As a result, the number of inter-chiplet gates should represent a strong proxy metric for the quality of a compiler solution, and thus minimizing the number of inter-chiplet gates should result in higher-fidelity, faster-executing quantum circuits. Accordingly, SEQC consistently produces results that have significantly fewer inter-chiplet gates compared to Qiskit-based compilers (Figure 7). Specifically, compilation with SEQC-STA results in $82.43\%$ fewer inter-chiplet gates on average compared to Qiskit (up to $92.94\%$). Compared to PA-Qiskit, compilation with SEQC-STA results in $64.92\%$ fewer inter-chiplet gates on average, up to $78.62\%$. It is worth mentioning that the inter-chiplet gates in the universal backend would be CZ gates, while the non-universal backend would be SWAP gates. For every inter-chiplet (CZ) gate in Qiskit, PA-Qiskit should have 2 inter-chiplet (SWAP) gates. However, we observe PA-Qiskit produces circuits with a roughly comparable number of inter-chiplet gates to Qiskit. In fact, as size increases, PA-Qiskit tends to result in fewer inter-chiplet gates than Qiskit. This outcome may indicate the presence of optimizations that are removing extraneous SWAPs.
@@
Figure 8. Relative gate count of SEQC normalized to the Qiskit baseline for devices (lower is better).
Figure 9. Relative circuit depth of SEQC normalized to the Qiskit baseline for devices (lower is better).
Gate count and circuit depth are static, backend-agnostic metrics that can serve as indicators for a quantum circuit’s expected fidelity and/or execution time, where circuit depth is the critical path of a quantum circuit in terms of individual gates (mike_and_ike). SABRE (zou2024lightsabrelightweightenhancedsabre; sabre_swap), moreover, explicitly minimizes gate count by incorporating it into its cost functions. We observe that SEQC typically produces results with higher gate count and circuit depth compared to Qiskit and PA-Qiskit (Figures 8 and 9). This is due to SEQC prioritizing the reduction of the number of inter-chiplet gates over total gate count or circuit depth. Nevertheless, SEQC-STA, in particular, is able to demonstrate competitive results in these metrics for larger circuits. In terms of relative circuit depth, SEQC-STA has comparable or slightly better results compared to PA-Qiskit at $500$ qubits (50 chiplets) and above. In terms of relative gate count, SEQC-STA has comparable or slightly better results compared to PA-Qiskit at $400$ qubits (40 chiplets) and above. For the $810$ and $1000$ qubit ($81$ and $100$ chiplet) trials, it achieves $14.11\%$ and $8.87\%$ gate count reduction over Qiskit, respectively, despite the latter using universal inter-chiplet links.
Fidelity
Figure 10. Relative estimated fidelity (ESP) of our proposed compilers normalized to the Qiskit baseline (higher is better).
Practical applications of quantum computing require some guarantee of program output correctness (measurements). Thus, the highest priority for quantum circuit compilation is to maximize fidelity. Fidelity is impacted by the decoherence constraints of individual qubits, as well as the errors on a gate-by-gate basis (mike_and_ike). Due to hardware limitations, we were unable to estimate fidelity though simulation. Instead, we leverage estimated success probability (ESP) (esp) to approximate fidelity, which can be computed statically. Specifically, we track the per-qubit ESP of a quantum circuit and report an averaged ESP for each benchmark from the geometric mean of the circuit measurements. Figure 10 presents the relative estimated fidelity of each of our compilers compared to the Qiskit compiler with respect to circuit size. There is a noticeable benefit in fidelity associated with the SEQC compiler, particularly with SEQC-STA. We observe an average improvement of $22.13\%$ over PA-Qiskit, up to a $62.87\%$ improvement. Moreover, it is important to mention that the Qiskit trial was ran on a backend with universal inter-chiplet links, which should translate to higher fidelity solutions. Nevertheless, SEQC-STA demonstrates an average improvement of $9.30\%$, up to a $49.99\%$ improvement.
@@
Figure 11. Estimated per-shot execution time of circuits produced by SEQC normalized to circuits produced by the Qiskit baseline (lower is better).
The per-shot execution time of a quantum circuit is determined by the critical path of gate operations weighted by gate duration. In addition to performance concerns, execution time indirectly affects fidelity on NISQ-era devices due to decoherence. In other words, it is imperative that quantum circuits execute within the decoherence time to ensure high-fidelity outcomes. Figure 11 presents the estimated per-shot execution time of our proposed compilers relative to the reference as a function of the number of qubits per circuit. As explained previously, we approximate the duration of an inter-chiplet SWAP as 4×4$\times$4 × the duration of an intra-chiplet SWAP (smith_2022_chiplets). Under this model, we observe that the SEQC compilers consistently outperform PA-Qiskit in execution by $24.19\%$ on average ($63.98\%$ max). It is notable that, despite the gate set affordances given to the Qiskit compiler, SEQC-STA is able to achieve parity in execution time starting from $600$ qubits ($60$ chiplets). For $810$ qubits ($81$ chiplets) and above, SEQC-STA actually compiles faster-executing quantum circuits.
@@
Figure 12. Compilation time of our evaluated compilers (lower is better).
Figure 13. Relative recurring compilation time of our proposed compilers normalized to the Qiskit baseline (lower is better).
@@
Near-term quantum devices feature rapidly fluctuating error characteristics that can change hourly, or even on a run-to-run basis (Tannu_2019_ErrorVariability; Murali_2019_NoiseMapping; dasgupta2021stability). Thus, across the entire execution lifetime of an application (quantum circuit), regular re-compilation is necessary to generate the highest-quality circuits with the most up-to-date characteristics possible for every execution. Traditional quantum compilers, like Qiskit and PA-Qiskit, approach each compilation tabula rasa, where each new compilation restarts from the original algorithmic circuit. It is challenging to speed up repeated recompilations without fundamental changes to their design, as noise-aware routing and synthesis algorithms, which leverage updated error characteristics, usually function from a device-agnostic starting point. In contrast, SEQC’s stratification stage enables stateful retention of previous compilations. It is agnostic to the ever-changing error characteristics of a quantum device, and instead only depends on the device’s static topology. Therefore, a stratified quantum circuit can be perennially reused over multiple rounds of compilation, and thus SEQC only repeats the elaboration stage.
@@
Nevertheless, although stratification is a one-time cost for the SEQC compiler, it must be scalable in order to be practical. In Figure 12, we plot the non-recurring compilation time of our evaluated compilers. For our SEQC compilers, we plot stratification time and stratification + elaboration time. These can be compared to the total solve time for the Qiskit and PA-Qiskit compilers. We demonstrate that all compilation times obey a quadratic trajectory ($R^{2}\geq 0.9928$).

### v4

@@ 4.2. Simulated Quantum Device Specifications
(a) Small-scale heavy-hex modules, on grid lattice.
(b) Large-scale modules with IBM Nighthawk-like topology.
Figure 6. Modeled topologies.
@@
CZ (intra-module) (backend_specs) $34.0$ $0.605\%$
CZ (inter-module) $126.0$ $2.420\%$
SWAP (inter-module) $702.4$ $10.23\%$
@@
Table 1. Specifications of Simulated Modular Backend.
We model two types of simulated devices differing in their inter-module topology and module size. To isolate the impact of architecture modularity, we evaluated on symmetrical modules based on a transmon qubit technology. One variant, denoted the heavy-hex backend (Figure 6(a)), is generated to conform to the modular architecture proposed in Smith et al. (smith_2022_chiplets), and features a 10-qubit, heavy-hexagon topology (heavy_hex). The other, denoted the grid backend (Figure 6(b)), uses modules with a 120-qubit grid topology, based on the upcoming IBM Nighthawk processor (IBM_2025_FTQC; Nighthawk_topology). Both simulated devices use a grid lattice topology of symmetric modules. The hardware specifications of the qubits and intra-module connections (detailed in Table 1) are sourced from Acharya et al. (backend_specs). To generate the expected fidelity of the inter-module SWAPs, we perform a simulated random benchmarking trial and calculate the error rate of a SWAP gate with these backend specifications. Following Smith et al. (smith_2022_chiplets), we conservatively model inter-module SWAPs with $4\times$ the error rate and $4\times$ the duration of an intra-module SWAP. Finally, note that for each valid module count, there may exist multiple valid module topologies. In this work, we generate our backends to use the “most-square” topology for a given module count. For example, a $12$-module machine would be a $3\times 4$ grid.
@@ 5. Experimental Results
We evaluate the Qiskit reference compiler, the peephole-augmented Qiskit (PA-Qiskit) compiler, SEQC using simulated annealing (SEQC-AN), and SEQC using STA (SEQC-STA) on metrics relating to the quality of the compiled circuits, on both the 10-qubit heavy-hex backend, and the 120-qubit grid backend. The Qiskit-based and SEQC compilers prioritize different heuristics in compilation. We examine how well each compiler is able to minimize its preferred heuristics, and how well that translates to practical solution quality. To demonstrate scalability, we analyze compilation time, specifically the recurring and non-recurring compilation times for repeating compilations of a single application, as well as memory requirements for the compilation. Relative metrics are normalized over the Qiskit reference compiler.
Number of inter-module gates
@@
The primary heuristic we minimize in SEQC is the number of inter-module gates in a quantum circuit. This is because we expect that the inter-module gates of a modular architecture contribute the most to error and execution time relative to other gates (at least by a factor of $4\times$ (smith_2022_chiplets)). As a result, the number of inter-module gates should represent a strong proxy metric for the quality of a compiler solution, and thus minimizing the number of inter-module gates should result in higher-fidelity, faster-executing quantum circuits. Accordingly, SEQC consistently produces results that have significantly fewer inter-module gates compared to Qiskit-based compilers (Figures 7(a) and 8(a)). For example, compilation with SEQC-STA results in $82.43\%$ fewer inter-module gates on average compared to Qiskit (up to $92.94\%$). on the heavy-hex backend. Meanwhile, compilation with SEQC-STA results in $96.31\%$ fewer inter-module gates on average compared to Qiskit (up to $98.32\%$) on the grid backend. Compared to PA-Qiskit, compilation with SEQC-STA results in $64.92\%$ fewer inter-module gates on average, up to $78.62\%$, on the heavy-hex backend. Compared to PA-Qiskit, compilation with SEQC-STA results in $94.31\%$ fewer inter-module gates on average, up to $97.83\%$ on the grid backend.
It is worth mentioning that the inter-module gates in the universal backend would be CZ gates, while the non-universal backend would be SWAP gates. For every inter-module (CZ) gate in Qiskit, PA-Qiskit should have 2 inter-module (SWAP) gates. However, we observe PA-Qiskit produces circuits with a roughly comparable number of inter-module gates to Qiskit. In fact, as size increases, PA-Qiskit tends to result in fewer inter-module gates than Qiskit. This outcome may indicate the presence of optimizations that are removing extraneous SWAPs.
(a) Relative inter-module gate count
(lower is better).
(b) Relative estimated fidelity (ESP)
(higher is better).
Figure 7. Main circuit performance metrics: inter-module gate count and estimated fidelity, 10-qubit heavy-hex backend.
(a) Relative inter-module gate count
(lower is better).
(b) Relative estimated fidelity (ESP)
(higher is better).
Figure 8. Main circuit performance metrics: inter-module gate count and estimated fidelity, 120-qubit grid backend.
Fidelity
Practical applications of quantum computing require some guarantee of program output correctness (measurements). Thus, the highest priority for quantum circuit compilation is to maximize fidelity. We estimate fidelity through estimated success probability (ESP) (esp). Specifically, we track the per-qubit ESP of a quantum circuit and report an averaged ESP for each benchmark from the geometric mean of the circuit measurements. Figures 7(b) and 8(b) presents the relative estimated fidelity of each compiler compared to the Qiskit compiler with respect to circuit size. There is a noticeable benefit in fidelity associated with the SEQC compiler, particularly with SEQC-STA. On the heavy-hex backend, we observe an average improvement of $22.13\%$ over PA-Qiskit, and up to $62.87\%$. On the grid backend, SEQC-compiled circuits achieve $3.51\%$ higher fidelity on average ($22.50\%$ max) over PA-Qiskit.
SEQC similarly outperforms Qiskit by a wide margin. It is important to emphasize here that SEQC outperforms Qiskit even when the latter is afforded significant allowances (no gate restrictions, and much lower inter-module gate errors). The Qiskit trial backend supports universal inter-module links, in contrast to the SEQC backend, and with only a $2.42\%$ error for inter-module CZs, compared to $10.23\%$ error for SEQC’s inter-module SWAPs. All these allowances should translate to higher fidelity solutions for Qiskit. Nevertheless, even though much more constrained, SEQC-STA demonstrates an average improvement of $9.30\%$, up to a $49.99\%$, on the heavy-hex backend, and $32.30\%$ ($63.36\%$ max) on the grid backend.
(a) Relative total gate count.
(b) Relative circuit depth.
(c) Relative per-shot execution time.
Figure 9. Total gate count, circuit depth, and per-shot execution time, 10-qubit heavy-hex backend (lower is better).
(a) Relative total gate count.
(b) Relative circuit depth.
(c) Relative per-shot execution time.
Figure 10. Total gate count, circuit depth, and per-shot execution time, 120-qubit grid backend (lower is better).
@@
Gate count and circuit depth are static, backend-agnostic metrics that can serve as indicators for a quantum circuit’s expected fidelity and/or execution time, where circuit depth is the critical path of a quantum circuit in terms of individual gates (mike_and_ike). SABRE (zou2024lightsabrelightweightenhancedsabre; sabre_swap), moreover, explicitly minimizes gate count by incorporating it into its cost functions. We observe that SEQC typically produces results with higher gate count (Figures 9(a) and 10(a)) and circuit depth (Figures 9(b) and 10(b)) compared to Qiskit and PA-Qiskit. This is due to SEQC prioritizing the reduction of the number of inter-module gates over total gate count or circuit depth. Nevertheless, SEQC-STA, in particular, is able to demonstrate competitive results in these metrics for larger circuits.
SEQC-STA has comparable or slightly better results compared to PA-Qiskit at $500$ qubits (50 modules) and above for both metrics on the heavy-hex backend. For the $810$ and $1000$ qubit ($81$ and $100$ module) trials on the heavy-hex backend, SEQC-STA achieves $14.11\%$ and $8.87\%$ gate count reduction over Qiskit, respectively, despite the latter using universal inter-module links. For the grid backend, we observe similar behavior to the heavy-hex backend, where SEQC produces increasingly competitive results as the circuit size increases. However, we did not perform any trails large enough to witness the point where SEQC-STA would begin to produce better results than the Qiskit baselines
@@
The per-shot execution time of a quantum circuit is determined by the critical path of gate operations weighted by gate duration. In addition to performance concerns, execution time indirectly affects fidelity on NISQ-era devices due to decoherence. In other words, it is imperative that quantum circuits execute within the decoherence time to ensure high-fidelity outcomes. Figures 9(c) and 10(c) presents the estimated per-shot execution time of our proposed compilers relative to the reference as a function of the number of qubits per circuit. As explained previously, we approximate the duration of an inter-module SWAP as $4\times$ the duration of an intra-module SWAP (smith_2022_chiplets). Under this model, we observe that the SEQC compilers consistently outperform PA-Qiskit in execution by $24.19\%$ on average ($63.98\%$ max) on the heavy-hex backend. It is notable that, despite the gate set affordances given to the Qiskit compiler, SEQC-STA achieves parity in execution time starting from $600$ qubits ($60$ modules). For $810$ qubits ($81$ modules) and above, SEQC-STA actually compiles faster-executing quantum circuits. Meanwhile, on the grid backend, SEQC-STA-compiled circuits execute up to $151.29\%$ faster over PA-Qiskit. Circuit execution time for SEQC-STA-compiled circuits is also faster than the Qiskit baseline from $480$ qubits ($4$ modules) onwards by $28.04\%$ on average, up to $54.82\%$, with SEQC-SA trailing close behind.
@@
Quantum devices feature rapidly fluctuating error characteristics that can change hourly, or even on a run-to-run basis (Tannu_2019_ErrorVariability; Murali_2019_NoiseMapping; dasgupta2021stability). Thus, across the entire execution lifetime of an application (quantum circuit), regular re-compilation is necessary to generate the highest-quality circuits with the most up-to-date characteristics possible for every execution. Traditional quantum compilers, like Qiskit and PA-Qiskit, approach compilation tabula rasa, where each new compilation restarts from the original algorithmic circuit. It is challenging to speed up repeated recompilations without fundamental changes to their design, as noise-aware routing and synthesis algorithms, which leverage updated error characteristics, usually function from a device-agnostic starting point. In contrast, SEQC’s stratification stage enables stateful retention of previous compilations. It is agnostic to the ever-changing error characteristics of a quantum device, and instead only depends on the device’s static topology. Therefore, a stratified quantum circuit can be perennially reused over multiple rounds of compilation, and thus SEQC only repeats the elaboration stage.
@@
Nevertheless, although stratification is a one-time cost for the SEQC compiler, it must be scalable in order to be practical. In Figures 11(a) and 12(a), we plot the non-recurring compilation time of our evaluated compilers. For our SEQC compilers, we plot stratification time and stratification + elaboration time. These can be compared to the total solve time for the Qiskit and PA-Qiskit compilers. We demonstrate that all compilation times obey a quadratic trajectory ($R^{2}\geq 0.9883$).